OpenStudy (anonymous):

Given that f(x) = x2 + 2x + 3 and g(x) = quantity of x plus four, over three, solve for f(g(x)) when x = 2

OpenStudy (campbell_st):

well you need to subsitute \[\frac{x + 4}{3}\] into f(x)

OpenStudy (anonymous):

i got 11, but i'm not sure i read your equations right.

OpenStudy (campbell_st):

then when you have substituted, and simplified replace x with 2 then evaluate

OpenStudy (anonymous):

11 is a option i think ill give it a try. but thank you.

OpenStudy (campbell_st):

so \[f(\frac{x +4}{3}) = (\frac{x +4}{3})^2 + 2(\frac{x +4}{3}) + 3\]

OpenStudy (campbell_st):

so next replace x with 2 and evaluate

OpenStudy (anonymous):

Solution. I Think this is the answer.. f(x) = x2 + 2x + 3 g(x) = x+4/3 Now we can get the f(g(x)) using this equation.. f(x+4/3) = (x+4/3)2 + 2(x+4/3) + 3 then x = 2 f(x+4/3) = (2+4/3)2 + 2(2+4/3) + 3 f(x+4/3) = 4 + 4 + 3 f(x+4/3) = 11 I think this is the answer..

OpenStudy (whpalmer4):

@MaxLure while you got the right answer, it is very sloppy to write \(\dfrac{x+4}{3}\) as \(x+4/3\) as by the rules of operator precedence, they are not equivalent. You should write \((x+4)/3 \) if you insist on forgoing the formatting tools the site provides.

OpenStudy (anonymous):

hahah sorry .. i don't know how to use this tools over here..